Question 9 10 pts Let A-(3.1i-2.33+1.9k) cm and B have components B, -2.7 cm, B,-1.5 cm, with no y component. What is A+B? O 1.21 cm (0.4i - 2.33+3.4k) cm CTL O 7.39 cm 3.9 cm o (5.8i -2.33+0.4k) cm Question 10 10 pts What is the magnitude (in crn of the vector (-101 + 2 8 . 2.3 k) cm? (Here k is the unit vector in the z direction .. there does not appear to be an easy way to make k hat in html.) Hint: the Pythagorean theorem works in three dimensions as well as two dimensions. both questions
Answer #1

Given A (3.1i _ 2.3j +1.9k)cm and B = (2.7i 03-1.5k)cm

vec{A}+vec{B}=?

vec{A}+vec{B}=(3.1hat{i} -2.3hat{j}+1.9hat{k}) +(2.7hat{i} +0hat{j}-1.5hat{k})

vec{A}+vec{B}=(3.1+2.7)hat{i} +(-2.3+0)hat{j}+(1.9-1.5)hat{k}

vec{A}+vec{B}=5.8hat{i} -2.3hat{j}+0.4hat{k}

ANSWER: Option 5

================

Magnitude of 1.02.8i- 2.3k

R = V/(-1,0)2 + (2.8)2 + (-2.3)2

ANSWER: R-3.759cm

=======================